[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Mar 28, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Mar 28 12:07:44 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark DeMaria

----------

TAFB:
There was a period of GOES-east outage last night due to an anomaly on the  
satellite. Data was restored by about 3 UTC.
HSU:

Other items:
We are upgrading our AWIPS systems to 16.1.4 this morning. OPC could not do  
the service back up due to unexpected SL, but the AWIPS support staff  
assisting us said they could bring in extra help and get it done in less  
than two hours. We do have HFO back up so we decided to go ahead with the  
upgrade. In a worst case scenario, the Miami WFO said we could use their  
AWIPS to submit the required TAFB text products if the upgrade takes longer  
than expected. If we use the WFO, the text products would be generated  
manually rather than from the GFE grids.
NHC RFCs:
RFC 164 will be implemented today to move the processing of TAFB animated  
gifs off Mike Brennan's workstation and onto the operational compute  
servers supported by TSB.

RFC 163 to bring down the experimental GOES-R QPE product to N-AWIPS will  
be implemented on Tuesday.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160328/9e29fa79/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list